# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

### v1

@@ Abstract
Document understanding is a challenging task to process and comprehend large amounts of textual and visual information. Recent advances in Large Language Models (LLMs) have significantly improved the performance of this task. However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially in academic papers. In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) which is designed to enhance multimodal question-answering (QA) for long PDF documents. PDF-WuKong incorporates a sparse sampler that operates on both text and image representations, significantly improving the efficiency and capability of the MLLM. The sparse sampler is integrated with the MLLM’s image encoder and selects the paragraphs or diagrams most pertinent to user queries for processing by the language model. To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of academic papers sourced from arXiv, multiple strategies are proposed to generate automatically $1M$ QA pairs along with their corresponding evidence sources. Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal PDF understanding, surpassing proprietary products by an average of 8.6% on F1. Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
@@ 1 Introduction
To handle lengthy documents, current large models mainly fall their input into two separate modalities. One is the plain text modality. As shown in Fig. 1 (a), they focus only on the parsed pure text content and lack multimodal understanding of visual elements such as charts and figures within the documents. They usually rely on long-context LLMs [3, 4, 5], whose efficiency and accuracy decrease as the length of the document increases. Retrieval Augmented Generation (RAG) [6, 7, 8, 9] is also the main auxiliary means of these LLMs, but it requires the introduction of additional models and lacks the exploration of multimodal RAG.
Figure 1: Method comparison for long multi-page PDF document understanding. (a) Plain text solution: long-context/RAG LLMs for parsed pure text content. (b) Purely visual solution: VDU models for page-level encoding and feature interaction. (c) Our method is based on end-to-end sparse sampling for long PDFs with interleaved text and images.
Another approach is to input documents in a purely visual modality, treating each page as an image. However, most current Visual Document Understanding (VDU) models models [10, 11, 12] only focus on visual QA on single-page documents. They usually rely on larger input resolution, resulting in more tokens, and thus have difficulty processing multi-page long documents. Recently, some multi-page VDU models can understand 8-page [13] or 20-page [14] documents. They usually encode each page separately and then perform page-level visual feature interactions such as concatenation [15, 16], as shown in Fig. 1 (b). However, more pages generate more visual tokens, which brings greater resource consumption to LLM. This makes them inefficient and unable to process longer documents.
@@
In order to simultaneously represent and understand the multimodal content of documents and further improve the ability to process long PDF documents, we construct a training dataset specifically for academic paper PDFs. The academic paper PDF is a kind of typical document that contains rich interleaved text and images, which can intuitively reflect the challenges of our task and the advantages of our model. The dataset contains complete PDF documents, professional academic questions, answers, and evidence sources for the answers, based on multiple construction strategies. We also provide a corresponding benchmark named PaperPDF.
The experimental results substantiate the effectiveness and efficiency of our approach on the task of long multimodal PDF understanding. PDF-WuKong significantly outperforms potential open-source models that may be applied to this task. It also surpasses some proprietary products for document understanding on our proposed PaperPDF benchmark. As the number of document pages increases, its accuracy and efficiency will not decrease significantly. It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE [17]. Besides, for the recent benchmark MM-NIAH [18] of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters. Our model achieves the best performance on multimodal content with a context length of 64K.
@@
  • We propose a PDF multimodal question answering dataset (PaperPDF) with $1M$ QA pairs for training and $6k$ QA pairs for evaluation.
@@
Purely visual Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], DocPedia [24], XC2-4KHD [25] and InternVL-V1.5 [11]
Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [26]
@@ 2.1 Document Understanding Datasets
Earlier document understanding datasets only focused on the NLP tasks such as summarization [27], and QA [28] of plain text. Meanwhile, there were several visual document datasets mainly aimed at text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34]. Recently, more datasets have been proposed for multimodal document QA across various scenarios. For example, DocVQA [35] and OCRVQA  [36] provide single-page QA data on books and business documents. ChartQA [37] and ChartX [38] focus on the visual reasoning for chart documents. ArXivQA [39] extracts scientific figure-caption data from the arxiv papers to enhance the academic ability of MLLMs. InfoVQA [40] contains many infographic documents which are a combination of textual, graphical and visual elements. However, these visual document datasets only define the single-page VQA task, and current MLLMs [41, 11] have achieved remarkable performance on it.
There are also some multi-page QA datasets [15, 17, 14] that require the model to understand the content relationship via multi-hop reasoning and capture the crucial information from multi-page documents. MP-DocVQA [14] extends DocVQA [35] by adding the context pages. DUDE [17] constructs multi-page QA data from multi-industry and multi-domain documents. Recently, DocGenome [42] was constructed as a scientific document benchmark. MM-NIAH [18] is a benchmark evaluating the capability of MLLMs to comprehend long multimodal documents, which requires the model to answer according to the key information scattered throughout the document. However, the answers in these datasets lack evidence and cannot provide reliable interpretability, especially for questions that require referring to multiple pieces of evidence in long documents.
@@ 2.2 Document Understanding Methods
Existing document understanding methods typically focus on either plain text or a limited number of document images. Methods that rely on pure text modality aim to process documents by first converting them into plain text through document parsing or Optical Character Recognition (OCR). They then employ efficient long-context mechanisms to handle long texts, such as sparse attention [4], memory networks [5], or position interpolation [3]. Besides, the methods based on retrieval-augmented generation (RAG) [6, 19, 9] also show impressive capabilities for long texts. While these approaches can integrate visual elements by including image captions or transforming images into natural language descriptions, they struggle with fine-grained understanding of visual information. This restricts their effectiveness in tasks requiring detailed interpretation of textual and visual components within documents.
Another solution is visual document understanding with purely visual input, treating each document page as an image. Many MLLMs such as UniDoc [20], mPLUG-DocOwl [21] and Vary [12] can perform this task in an OCR-free manner. Vary [12] employs an extra SAM-style [43] vision vocabulary specific to document and chart data, enabling direct encoding of entire pages with high compression ratios. Other researchers have advanced the understanding of high-resolution document pages by dividing input images into smaller patches, such as UReader [22], TextMonkey [10], and LLaVA-NeXT [23]. DocPedia [24] processes high-resolution images in the frequency domain via the DCT transformation. InternLM-XComposer2-4KHD [25] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration to capture more details. The reliance on high resolution results in a higher number of tokens and cannot be extended to multi-page documents.
Besides, there are several models specifically designed for multi-page documents. Hi-VT5 [14] and GRAM [16] are two professional models for multi-page QA. They combine image tokens of multiple pages through hierarchical transformer architecture [14] or global-local reasoning [16], being able to handle up to 20 pages. Fox [13] unifies all image tokens of up to 8 pages into a sequence to achieve multi-page QA. mPLUG-DocOwl2 [26] compress each high-resolution document image into 324 tokens, guided by low-resolution global visual features. These models encode each page separately and then perform page-level visual feature interactions. However, more pages generate more visual tokens, which brings greater resource consumption to LLM. This makes them inefficient and unable to process longer documents. Given this, we propose first to parse the lengthy document into interleaved text and image content and then perform sparse sampling in an end-to-end manner. Tab. 1 summarizes the attributes of each type of method.
@@ 3.1 Overview
In order to achieve multimodal understanding of long PDF documents and alleviate the drawbacks of existing models that treat PDF documents as a single modality of plain text or images, we propose PDF-WuKong. The core motivation is that users’ queries are often only related to a small number of text blocks or diagrams in a long document. Therefore, to improve the accuracy and efficiency of MLLM, we design a sparse sampler integrated with a multimodal large language model in an end-to-end manner.
@@
Specifically, our pipeline consists of three parts: a document parser, a sparse sampler and a large language model, as shown in Fig. 2. The document parsing stage first converts the input PDF document into machine-readable content of interleaved text and images, Then, the sparse sampler encodes the text blocks and images separately and caches their embeddings. When a user inputs a query, the most relevant content can be sampled using a simple similarity measure. Finally, the query and the sampled tokens are input into the LLM to generate the answer. Algorithm 1 shows the detailed steps of this method.

### v2

@@ Abstract
Multimodal document understanding is a challenging task to process and comprehend large amounts of textual and visual information. Recent advances in Large Language Models (LLMs) have significantly improved the performance of this task. However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially for academic papers. In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) which is designed to enhance multimodal question-answering (QA) for long PDF documents. PDF-WuKong incorporates a sparse sampler that operates on both text and image representations, significantly improving the efficiency and capability of the MLLM. The sparse sampler is integrated with the MLLM’s image encoder and selects the paragraphs or diagrams most pertinent to user queries for processing by the language model. To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of English and Chinese academic papers. Multiple strategies are proposed to automatically generate 1.1 million QA pairs along with their corresponding evidence sources. Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal document understanding, surpassing proprietary products by an average of 8.6% on F1. Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
@@ 1 Introduction
To handle lengthy documents, current research in multimodal document understanding with LLMs primarily follows two mainstream technical routes. The first route is based on pure text modality understanding. As shown in Fig. 1(a), these approaches typically consider the parsed OCR results from PDF documents, and convert all visual elements into textual representations (e.g., captions and OCR content extracted from figures and tables). They then employ long-context LLMs [3, 4, 5] or utilize Retrieval Augmented Generation (RAG) techniques [6, 7, 8, 9] to process the textual content. The main disadvantage of this approach is the significant loss of visual information inherent in multimodal documents, making it challenging to support answering fine-grained visual-related questions.
Figure 1: Method comparison for long multi-page PDF document understanding. (a) Plain text solution: long-context/RAG LLMs for parsed pure text content. (b) Pure vision solution: VDU models for page-level encoding and feature interaction. (c) Our method is based on end-to-end sparse sampling for long PDFs with interleaved text and images.
The second route is based on pure visual modality understanding. As illustrated in Fig.1(b), this approach generally avoids parsing PDF documents and instead treats each page as an image, and utilizing multimodal LLMs for document understanding tasks[10, 11, 12]. However, high-resolution images and numerous pages generate a large number of visual tokens, significantly increasing the models’ input token length and leading to scalability issues. This makes it challenging to efficiently process multi-page long documents. While some recent multi-page VDU models handle up to 8 pages [13] or 20 pages [14], they typically encode each page separately and perform page-level visual feature interactions such as concatenation [15, 16]. Nevertheless, as the number of pages grows, computational resource consumption still escalates substantially, rendering these models inefficient for processing longer documents.
@@
In order to simultaneously represent and understand the multimodal content of documents and further improve the ability to process long PDF documents, we construct a training dataset specifically for English and Chinese academic paper PDFs. The academic paper PDF is a kind of typical document that contains rich interleaved text and images, which can intuitively reflect the challenges of our task and the advantages of our model. The dataset contains complete PDF documents, professional academic questions, answers, and evidence sources for the answers, based on multiple construction strategies. We also provide a corresponding bilingual benchmark named PaperPDF.
We train PDF-WuKong on PaperPDF dataset, complemented by general-domain document question-answering datasets. Experimental results substantiate the effectiveness and efficiency of our approach to the task of long multimodal PDF understanding. PDF-WuKong significantly outperforms potential open-source models that may be applied to this task. It also surpasses some proprietary products for document understanding on our proposed PaperPDF benchmark. As the number of document pages increases, its accuracy and efficiency will not decrease significantly. It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE [17]. Besides, for the recent benchmark MM-NIAH [18] of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters. Our model achieves the best performance on multimodal content with a context length of 64K.
@@
  • We propose a bilingual PDF multimodal question answering dataset (PaperPDF) with $1.1M$ QA pairs for training and $10k$ QA pairs for evaluation.
@@
Pure vision Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], XC2-4KHD [24], InternVL-V1.5 [11]
Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [25], CREAM [26]
@@ 2.1 Document Understanding Datasets
Early datasets focused on NLP tasks like summarization [27] and QA [28] of plain text, while visual document datasets targeted text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34]. Recent multimodal document QA datasets include DocVQA [35] and OCRVQA [36] for single-page documents, ChartQA [37] and ChartX [38] focusing on visual reasoning in charts. Datasets like ArXivQA [39] and InfoVQA [40] enhance MLLMs’ abilities on academic and infographic documents. However, these datasets are limited to single-page tasks, and current MLLMs [41, 11] perform well on them.
Multi-page QA datasets like MP-DocVQA [14], DUDE [17], DocGenome [42], and MM-NIAH [18] require understanding content relationships via multi-hop reasoning. Yet, answers in these datasets lack evidence and reliable interpretability, especially for questions needing multiple pieces of evidence from long documents.
@@ 2.2 Document Understanding Methods
Existing methods focus on plain text or limited document images. Text-based approaches convert documents into plain text using OCR and then employ long-context mechanisms like sparse attention [4], memory networks [5], or position interpolation [3]. Retrieval-augmented generation methods [6, 19, 9] also handle long texts effectively. These approaches struggle with fine-grained visual understanding.
Another solution, visual document understanding, treats each page as an image. MLLMs like UniDoc [20], mPLUG-DocOwl [21], and Vary [12] perform OCR-free understanding. Models such as UReader [22] and TextMonkey [10] divide high-resolution pages into patches. InternLM-XC2-4KHD [24] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration. However, reliance on high resolution increases token counts and isn’t scalable to multi-page documents.
For multi-page documents, models like Hi-VT5 [14], GRAM [16], Fox [13], CREAM [26] and mPLUG-DocOwl2 [25] encode pages separately and perform page-level interactions. More pages generate more visual tokens, increasing resource consumption and inefficiency for longer documents. Thus, we propose parsing documents into interleaved text and images, followed by sparse sampling in an end-to-end manner. Tab. 1 summarizes these methods.
@@ 3.1 Overview
Our pipeline consists of three components: a document parser, a sparse sampler, and a large language model, as shown in Fig. 2. The document parsing stage converts input PDFs into machine-readable content with interleaved text and images. The sparse sampler then encodes and caches embeddings for text blocks and images separately. When receiving a user query, it retrieves the most relevant content through similarity matching. Finally, the query and sampled tokens are fed into the LLM for answer generation. The detailed procedure is outlined in Algorithm 1 in the appendix.
3.2 Document Parsing
3.3 Sparse Sampling
During the training, for the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query. Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec. 4). We randomly select two text blocks and two images from the remaining text blocks and images as negative samples. Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$. An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}}{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}+\displaystyle\sum_{e_{j}\in E_{N}}e^{\frac{\text{sim}(e_{q},e_{j})}{\tau}}},$ (1)
where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples. $\tau$ is the temperature parameter that controls the scale of the similarity scores. $P$ is the number of positive samples. By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
During the inference, we pre-encode all text blocks and images and cache all candidate embeddings. When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings. Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query. Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences. Moreover, the multimodal embedding cache further optimizes inference time. The pseudocodes for training and inference are shown in the appendix.
3.4 Answer Generation
At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency. Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence.
PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and QA.
4 PaperPDF Dataset
4.1 Overview
Our motivation for creating PaperPDF is threefold: (1) In long document QA contexts, answers often derive from specific segments, with other content acting as noise and complicating MLLM reasoning; (2) Existing datasets are either limited to single-page documents or lack fine-grained evidence ground truth, hampering the training of our sparse sampler; (3) There is currently a lack of a bilingual multimodal PDF QA dataset. Therefore, we introduce a method for automatically generating question-answer pairs from long documents and present PaperPDF, a dataset designed for both training and evaluation.
4.2 Dataset construction strategy
The PaperPDF dataset is constructed through a four-step process as shown in Fig. 3: document parsing, evidence extraction, QA generation, and data filtering. First, $100k$ PDFs are parsed to extract text and image chunks. Then we extract some evidence from each PDF according to our specific rules. Given the evidence and the generation prompt as the input, we use Gemini Pro [43] to generate questions and answers for the training set due to its free and rapid accessibility. GPT-4V [44] is used for the test set to ensure high evaluation quality. After obtaining the preliminary dataset, we conduct automatic filtering and remove abnormal data. For the test data, we further perform manual filtering from multiple aspects to ensure its quality. According to different evidence extraction rules, the triplets in PaperPDF can be divided into two categories.
Single-evidence Q-E-A triplets. The questions in these triplets can be answered based on a single text chunk or a diagram. Therefore, the evidence can be categorized into Text-only and Image-only. These triplets enable PDF-WuKong to initially acquire the capabilities of sparse sampling and long multimodal document understanding.
Multi-evidence Q-E-A triplets. These triplets require reasoning across multiple chunks, involving combinations of text and images. The types include Image-text (derived from a text chunk and associated images), Section (generated from all chunks in a section), and Cross-paragraph (involving related paragraphs across a document). For the third type, semantic summarizations for each paragraph are conducted first, followed by a selection of several related text chunks for QA generation. These multi-evidence triplets enhance our model’s multi-hop reasoning capability.
In total, we obtained $1.1M$ bilingual training data and $10k$ testing data. The dataset statistics are shown in Tab. 2. The appendix contains more data statistics.
Category Train (En/Zh) Test (En/Zh)
Single Text-only 249k/12k 2939/296
Image-only 21K/40k 212/1018
Multi Image-text 250k/53k 2566/2150
Section 499k/7k 255/394
Cross-paragraph 1.2k/0 118/0
Table 2: The statistics of PaperPDF in English and Chinese.
5 Experiments
5.1 Implementation Details
The LLM and the vision encoder are initialized from IXC2-VL-4KHD [24] and the maximum number of dynamic tiles is set as 16. The text encoder is initialized from BGE-M3 [45]. We train the model by leveraging several document datasets including PaperPDF, DocVQA [35], ChartQA [37], InfoVQA [40], MPDocVQA [46], and DUDE [17]. Before both training and testing, PaperPDF is parsed using Grobid [47] and MinerU [48], while the other datasets are processed following their default instructions. The training is conducted for one epoch using 128 Ascend 910B NPUs with a learning rate of 4e-5. We set the top 5 sampling results as input to the LLM.
For the model evaluations on the PaperPDF dataset, we use three objective metrics ANLS, F1, and Rouge to report the quantitative results on the full test set. To evaluate the semantic correctness of the answer, we introduce GPT-Acc to determine whether the output is correct. Considering the expensive cost of GPT-4 evaluation, we randomly selected two subsets with 50 English PDFs and 30 Chinese PDFs for GPT-Acc calculation. They contain 488 and 317 QA samples, respectively.
Model #Param English Chinese Token
ANLS F1 Rouge GPT-Acc ANLS F1 Rouge GPT-Acc
Plain Text Solution
IXC2-VL [49] 8B 27.4 30.8 32.8 37.6 21.1 28.5 28.7 30.8 4644
IXC2-VL-RAG [49] 8.5B 32.4 34.0 32.4 48.4 23.3 29.8 32.3 38.1 623
InternVL2 [11] 8B 19.5 28.0 27.6 37.5 24.4 28.2 27.7 30.8 4051
InternVL2-RAG [11] 8B 29.8 29.8 28.3 50.2 24.6 27.8 26.8 43.0 583
Pure Vision Solution
Hi-VT5 [46] 0.3B 13.5 3.1 3.7 15.2 - - - - 11589
IXC2-VL [49] 8B 27.1 24.8 25.1 20.5 15.9 19.5 22.2 27.4 4712
InternVL2 [11] 8B 29.4 33.1 35.0 35.5 20.3 33.9 37.8 37.7 5008
Parsed Interleaved Text and Images
IXC2-VL [49] 8B 27.8 31.2 32.6 37.7 22.5 29.5 29.2 31.4 6217
InternVL2 [11] 8B 33.4 36.2 36.6 54.3 28.5 40.6 42.0 54.7 6220
PDF-WuKong (ours) 8.5B 41.9 43.5 40.9 77.5 40.9 47.8 48.6 57.8 2107
Table 3: Performance comparison with open-source models for long PDF understanding on PaperPDF. The best results are marked bold and the second results are underlined.
Model ANLS F1 Rouge GPT-Acc
Gemini pro [50] 26.6 29.0 29.8 67.9
Kimi [51] 28.5 33.6 31.1 74.7
ChatGLM [52] 31.2 35.4 32.0 73.5
Qwen [53] 36.0 40.3 35.5 78.1
PDF-WuKong (ours) 41.8 43.2 40.7 77.5
Table 4: Performance comparison with commercial products (tested in Sep 2024) for long PDF understanding. The results are tested on a subset of 50 English PDFs. Note: These are the products based on the models rather than the models themselves.
5.2 Long PDF Understanding
To assess the effectiveness of our model in understanding long PDF documents, we conduct comprehensive experiments comparing it with both open-source models and commercial products on the PaperPDF dataset. Due to limited capabilities of traditional document understanding models  [46, 16, 54, 55, 56] and the huge resource costs in advanced MLLMs  [49, 24, 11], achieving deep understanding of lengthy PDF documents remains a highly challenging task. To handle this task, we explore three approaches to input the PDF documents into these MLLMs: pure text content, page images, and parsed interleaved text-image content. For some baselines with plain text input modality, we also report results based on retrieval enhancement techniques  [45]. The experimental results are reported in Tab. 3 and Tab. 4.
Several key conclusions can be drawn from the results. Firstly, inputting parsed interleaved text-image content generally outperforms multiple page images. The main reason is the limited input resolution and number of tokens that the model can accept. Secondly, when handling tokens of similar scale, inputting parsed interleaved text-image content yields better performance compared to pure text content. It is obvious that diagram information in the PDF plays a crucial role for document comprehension. This also indicates that PaperPDF places rigorous requirements on the visual information within documents. Additionally, we observe that for the InternVL2 model, the approach of parsing interleaved text-image content as input outperforms InternVL2-RAG. However, the opposite conclusion is drawn for IXC2-VL. We hypothesize that this discrepancy may be due to the max length setting in IXC2-VL, which could cause the model to overlook some critical information. Finally, benefiting from the inclusion of the spare sampler, our proposed PDF-WuKong model not only surpasses the existing state-of-the-art open-source model InternVL2 by approximately 7% on both the Chinese and English subsets, but also demonstrates competitive performance comparable to proprietary products. Moreover, due to the integration of the sparse sampler, PDF-WuKong maintains efficiency in inference token cost.
Model # param ANLS F1 ROUGE
Qwen-VL [57] 9.6B 26.4 19.6 18.3
Monkey [58] 9.8B 30.0 24.4 22.3
mPLUG-Owl2 [25] 8.2B 19.5 20.3 22.7
Emu2-Chat [59] 37B 26.0 24.4 23.4
MiniCPM-2.5 [60] 8.5B 31.8 28.2 24.8
IXC2-VL [49] 8B 23.4 20.8 21.3
IXC2-4KHD [24] 8B 24.5 20.0 18.0
CogVLM2 [61] 17B 24.8 27.4 26.3
[…]
Table 5: Performance comparison with other DocVLMs for PDF multimodal understanding on the Single-Evidence Subset. $\dagger$ denotes the page image input, aligning with other models in the table; * indicates that our model utilizes parsed content as input.
To compare with more open-source document MLLMs, considering that most of these models can only handle single-page documents, we construct a subset of the PaperPDF benchmark, only containing test samples with single evidence. Therefore, we provided all models with only one page containing the evidence as their input. The pages are input in the form of images. Our PDF-WuKong can accept input in two formats. One is the page image and another is the parsed page. As shown in Tab. 5, our model’s capability on this subset is significantly better than other document models. Moreover, the document parsing-based paradigm used in our approach is superior to the purely visual paradigm.
5.3 Document-oriented VQA
To validate the strong capability of our model in other document understanding scenarios, we conduct experiments on several public benchmarks and compare PDF-WuKong with other representative models. First, we evaluate the performance of PDF-WuKong on single-page document datasets [35, 37, 40]. As shown in Tab. 6, our model achieves comparable performance on single-page document understanding. This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question-answering tasks.
In addition, we assess the performance of traditional specialized models and MLLMs on two existing multi-page document QA datasets. The results shown in Tab. 7 indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document MLLM DocOwl2 [25]. Notably, on complex multi-page document datasets like DUDE [17], PDF-WuKong outperforms GPT-4V [44]. This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
Furthermore, we conduct zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18]. As shown in Tab. 8, our model uses the fewest parameters yet achieves the second-best performance. Although InternVL-V1-5-RAG surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model. Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, while that of other models significantly decreases. At a context length of 64K, PDF-WuKong even achieves the best performance, demonstrating its robustness in handling long-context multimodal inputs.
Model DocVQA ChartQA InfoVQA
Qwen-VL [57] 65.1 65.7 35.4
Monkey [58] 66.5 65.1 36.1
Text-Monkey [10] 73.0 66.9 28.6
MiniCPM-V-2.5 [60] 84.8 - -
Vary-base [12] 76.3 66.1 -
TextHawk [62] 76.4 66.6 50.6
IXC2-4KHD-16 [24] 84.9 80.1 60.8
DocOwl 2 [25] 80.7 70.0 46.4
CREAM [26] 79.4 - 53.6
PDF-WuKong (ours) 85.1 80.0 61.3
Table 6: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks.
Model MP-DocVQA DUDE
LayoutLMv3 [54] 55.1 20.3
Longformer [55] 55.1 27.1
BigBird [56] 58.5 26.3
Hi-VT5 [46] 61.8 35.7
DocFormerv2 [63] 76.4 48.4
GRAM [16] 83.0 53.4
GPT-4V (2024-06) [44] - 53.9
Idefics3-8B [64] 67.2 38.7
DocOwl2 [25] 69.4 46.7
CREAM [26] 65.3 52.5
PDF-WuKong (ours) 76.9 56.1
Table 7: Performance comparison with other DocVLMs for multi-page document understanding.
Model #param Overall 1K 4K 16K 64K
Emu2-Chat [59] 37B 8.8 38.9 18.2 0.0 0.0
VILA1.0-13b [65] 13B 15.7 41.9 33.2 8.6 0.1
llava-v1.6-13b [23] 13B 16.9 43.7 34.9 13.6 0.0
llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
InternVL1.5-RAG [11] 45B 46.1 59.5 50.1 44.9 39.3
PDF-WuKong (ours) 8.5B 43.3 53.0 43.9 43.0 42.1
Table 8: Performance comparison with other DocVLMs on MM-NIAH. The evaluation approach aligns with the benchmark.
5.4 Ablation Study
To comprehensively evaluate the effectiveness of our contributions, we conduct ablation studies focusing on the impact of the sparse sampler, the dataset, the document length, and sampling strategies. These experiments are based on English PaperPDF for training and evaluation to avoid interference from other factors.
Sparse sampler
To assess the effectiveness of the sparse sampler, we compared models trained with and without it. Without the sparse sampler, the MLLM struggled to process long documents with interleaved text and images, resulting in poor performance due to the large amount of irrelevant information. Introducing the sparse sampler significantly improved the model’s accuracy, as evidenced in Tab. 9, by efficiently selecting the most relevant content for each query. Furthermore, end-to-end joint training of the sparse sampler and the MLLM led to additional performance gains compared to training them separately. This indicates that our end-to-end optimization of multimodal representation and question answering can further promote the document understanding ability of MLLM.
Sparse Sampler End-to-End ANLS F1 ROUGE
11.1 5.1 5.0
40.3 42.3 39.8
42.6 43.6 40.2
Table 9: Ablation study on the impact of sparse sampler
Dataset
We retrain PDF-WuKong on various subsets of our English PaperPDF, and the results are shown in Tab. 10. Increasing the amount of training data leads to consistent improvements in the model’s accuracy, proving that our dataset follows scaling laws. In addition, we verify the effectiveness of our two data construction methods. Under the same data scale, multi-evidence data can enhance the model’s complex reasoning ability.
Dataset ANLS F1 ROUGE
100 k 38.7 40.1 37.5
500 k 41.6 43.5 40.8
1 M 42.6 43.6 40.2
Single (200k) 39.2 40.4 38.1
Multi (100k) + Single (100k) 40.0 41.6 38.9
Table 10: Ablation study on dataset setting
Document length
To understand the impact of document length on model performance and efficiency, we divide the test set into subsets based on the number of pages per document. Results in Fig. 4 demonstrate that our model’s performance and token counts remain relatively stable across documents of varying lengths. This stability indicates that the sparse sampler effectively reduces the input size to a reasonable level, regardless of the original document length. In contrast, the baseline MLLM without the sparse sampler is unable to handle long documents effectively. Its performance deteriorates significantly as the document length increases. The number of tokens also increased dramatically, resulting in huge resource consumption. These findings highlight the robustness of our model in processing long documents without sacrificing accuracy or incurring extra computational costs.
Figure 4: Ablation study of different document length
Sampling strategy
We explore the impact of different numbers of text blocks or diagrams selected by the sparse sampler. As shown in Tab. 11, setting a small top $k$ can lead to missing crucial information needed for accurately answering queries, thus reducing performance. Conversely, a larger top $k$ introduces redundant information and increases computational costs without significantly enhancing accuracy. To strike a balance between performance and resource efficiency, we use the top 5 as the default setting.
Sampling chunks ANLS F1 ROUGE Tokens
Top 1 39.20 38.28 35.26 1186
Top 3 42.09 43.06 39.69 1452
Top 5 42.59 43.63 40.19 1789
Top 10 43.01 44.22 40.67 2386
Top 15 43.19 44.57 42.08 2704
Top 20 43.42 45.02 42.30 3364
Table 11: Ablation study of different sampling strategy
6 Conclusion
We have presented PDF-WuKong, a novel MLLM that effectively addresses the challenges of understanding long PDF documents containing interleaved text and images. By introducing an end-to-end sparse sampling mechanism, our model efficiently extracts the most relevant paragraphs and diagrams in response to user queries, significantly reducing input token size and making the process independent of document length. We also constructed PaperPDF, a bilingual dataset with $1.1M$ question-answer pairs for training and $10k$ pairs for evaluation, specifically tailored for academic PDFs. Experimental results demonstrate that PDF-WuKong not only outperforms existing open-source models but also surpasses proprietary products by an average of 8.6% in F1 score on the long multimodal PDF understanding task. Our approach maintains high accuracy and efficiency even as document length increases, offering a scalable and interpretable solution for practical applications in document understanding.
Appendix A Algorithm
Algorithm 1 shows the detailed inference process of PDF-WuKong. The training pipeline is shown in Algorithm 2. Our PDF-WuKong can achieve efficient and accurate understanding of long PDFs with end-to-end sparse sampling.